### v2

@@ Abstract.
Let $G$ be a simple graph on $n$ vertices with chromatic number $\chi$, and let $\lambda_{n}$ denote the least adjacency eigenvalue. Solving a conjecture of Fan, Yu and Wang [Electron. J. Combin., 2012], we prove that when $3\leq\chi\leq n-1$, the chromatic number satisfies the following upper bound:
@@
We also compare this bound with the well known bound due to Wilf that $\chi\leq 1+\lambda_{1}$, where $\lambda_{1}$ denotes the largest eigenvalue. In particular we show that while Wilf’s bound is an upper bound for some parameters larger than $\chi$, this bound using $\lambda_{n}$ is not an upper bound for these parameters.
@@ 1. Introduction
In [3, Theorems 3.1 and 3.2], Fan, Yu and Wang established an upper bound for the chromatic number of a graph in terms of its order and least adjacency eigenvalue, which also appears in Stanić’s book [5, Theorem 3.11].
@@ 3. Main results
$G\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2}\right)=\left(K_{\frac{\chi}{2}}\cup\tfrac{\chi}{2}K_{1}\right)\vee\left(K_{\frac{n-\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right),$
@@ 4. A comparison with Wilf’s bound
As noted above, Wilf proved that $\chi(G)\leq 1+\lambda_{1}(G)$. In fact, as proved by Wu and Elphick [7, Lemma 3.2],
@@
When $\chi=2$, Theorem 3.1 simplifies to $\lambda_{n}\geq-n/2$, which was proved by Constantine [1] for all graphs. Theorem 3.2 is therefore true when $\chi=2$, and we can let $G=K_{3,3}$ for which $n=6$ and the least adjacency eigenvalue is $\lambda_{n}(G)=-3$. Plugging $\lambda=-3$ and $n=6$ into the right-hand side of (3.12) gives
$\left(\frac{n}{2}+1+\lambda\right)+\sqrt{\left(\frac{n}{2}+1+\lambda_{n}\right)^{2}-4(\lambda_{n}+1)\left(\lambda_{n}+\frac{n}{2}\right)}=2.$
@@
5. Concluding Remarks
In this paper, we resolved a conjecture proposed by Fan, Yu and Wang, proving that their upper bound on the chromatic number also holds for $\frac{n}{2}<\chi\leq n-1$. It is worth noting that they also proved the following result [3, Theorem 3.3]:

### v3

@@ Abstract.
Let $G$ be a simple graph on $n$ vertices and $m$ edges with chromatic number $\chi$, and let $\lambda_{n}$ denote the least adjacency eigenvalue. Solving a conjecture of Fan, Yu and Wang [Electron. J. Combin., 2012], we prove that when $3\leq\chi\leq n-1$, the chromatic number satisfies the following upper bound:
@@
We also compare this bound with the well-known bound due to Wilf that $\chi\leq 1+\lambda_{1}$, where $\lambda_{1}$ denotes the largest eigenvalue. In particular we show that while Wilf’s bound is an upper bound for some parameters larger than $\chi$, this bound using $\lambda_{n}$ is not an upper bound for these parameters. We conclude with a similar conjectured upper bound for $\chi(G)$, which uses $m$ in place of $n$.
@@ 1. Introduction
In [3, Theorems 3.1 and 3.2], Fan, Yu and Wang established an upper bound for the chromatic number of a graph in terms of its order and least adjacency eigenvalue, which also appears in Stanić’s book [6, Theorem 3.11].
@@ 3. Main results
$G\left(\tfrac{\chi}{2},\tfrac{n-\chi}{2}\right)=\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right)\vee\left(K_{\frac{\chi}{2}}\cup\tfrac{n-\chi}{2}K_{1}\right),$
@@ 4. A comparison with Wilf’s bound
As noted above, Wilf proved that $\chi(G)\leq 1+\lambda_{1}(G)$. In fact, as proved by Wu and Elphick [8, Lemma 3.2],
@@
When $\chi=2$, Theorem 3.1 simplifies to $\lambda_{n}\geq-n/2$, which was proved by Constantine [1] for all graphs. Theorem 3.2 is therefore true when $\chi=2$, and we can let $G=K_{3,3}$ for which $n=6$ and the least adjacency eigenvalue is $\lambda_{n}(G)=-3$. Plugging $\lambda_{n}=-3$ and $n=6$ into the right-hand side of (3.12) gives
$\left(\frac{n}{2}+1+\lambda_{n}\right)+\sqrt{\left(\frac{n}{2}+1+\lambda_{n}\right)^{2}-4(\lambda_{n}+1)\left(\lambda_{n}+\frac{n}{2}\right)}=2.$
@@
5. Extremal graphs
In addition to proving Theorem 1.1, Fan, Yu and Wang also proved the following result [3, Theorem 3.3]:
@@ 5. Extremal graphs
6. Concluding remarks
In this paper, we resolved a conjecture proposed by Fan, Yu and Wang, proving that their upper bound on the chromatic number also holds for $\frac{n}{2}<\chi\leq n-1$. Underpinning this bound is that for all graphs $\lambda_{n}\geq-n/2$. Powers [5] proved that $\lambda_{n}^{2}\leq m$, using the following simple proof:
$2m=\sum_{i=1}^{n}\lambda_{i}^{2}\geq\lambda_{1}^{2}+\lambda_{n}^{2}\geq 2\lambda_{n}^{2}.$
Wu and Elphick [8] proved that:
$\chi(\chi-1)\leq(\lambda_{1}+1)\lambda_{1}\leq 2m.$ (6.1)
By analogy with Theorem 1.1, this suggests the following conjecture, in which $m$ replaces $n/2$, $\chi(\chi-1)$ replaces $\chi$, and $\lambda_{n}^{2}$ replaces $\lambda_{n}$. Note that this conjectured bound is $\leq 2m$.
Conjecture 6.1.
For any non-empty graph $G$,
$\chi(\chi-1)\leq(m+1-\lambda_{n}^{2})+\sqrt{(m+1-\lambda_{n}^{2})^{2}-4(\lambda_{n}^{2}-1)(\lambda_{n}^{2}-m)}.$
Conjecture 6.1 is exact for $K_{n}$ and complete bipartite graphs, and the proof is trivial for bipartite graphs. We have verified Conjecture 6.1 for graphs of order at most 9, and graphs with order at most $100$ in the Wolfram Mathematica database. As for Theorem 3.2, the chromatic number in Conjecture 6.1 cannot be replaced by the list chromatic number or the coloring number.
This edge-based conjecture typically performs better than the vertex-based bound proved in this paper. For example for the connected graphs in the Mathematica database with 16 vertices:
  • Conjecture 6.1 outperforms the Fan-Yu-Wang bound (3.12) for 384 graphs;
  • the Fan-Yu-Wang bound (3.12) outperforms Conjecture 6.1 for 34 graphs; and
  • the bounds are equal for 2 graphs.